Question 4 Incorrect. Data on oxide thickness of semiconductors are as follows: 424 432 417 417 420 436 417 412 431 433 424 426 412 433 437 429 410 426 410 439 422 428 414 414 Consider this data as a sample of the population. (a) Calculate a point estimate of the mean oxide thickness for all wafers in the population. (Round your answer to 3 decimal places.) (b) Calculate a point estimate of the standard deviation of oxide thickness for all wafers in the population. (Round your answer to 2 decimal places.) (c) Calculate the standard error of the point estimate from part (a). (Round your answer to 2 decimal places.) (d) Calculate a point estimate of the median oxide thickness for all wafers in the population. (Express your answer to 1 decimal places.) (e) Calculate a point estimate of the proportion of wafers in the population that have oxide thickness greater than 430 angstrom. (Round your answer to 4 decimal places.)
Answer #1

putting above data in ascending order:

a)

point estimate of mean =423.458

b)

std deviaation =9.09

c)

std error =std deviation/sqrt(n)=1.86

d)

median =(n+1)/2 th value =12.5 th value =424

e)

proportion =7/24=0.2917